Predict the mode of decay of (a) plutonium-239, (b) Indium-120. Show your solution
1
Expert's answer
2021-07-14T05:48:50-0400

Part a

"plutonium-239 \implies \alpha - decay\\\\

_{94}^{239} Pa \to _{92}^{235} U + _{2}^{4} \alpha"

Part b

"Indium-120 \implies \alpha - decay\\\\

_{49}^{120} In \to _{50}^{120} Sn + _{-1}^{0} e"